I am having issues with determining a of n when the only given are two terms from an arithmetic series. The two terms given are a of 8= 75 and a of 20= 75. What would a of n be??
Answer by htmentor(1343)

I am having issues with determining a of n when the only given are two terms from an arithmetic series. The two terms given are a of 8= 75 and a of 20= 75. What would a of n be??
========================
The nth term of an arithmetic series can be written
a(n) = a + (n-1)d where a=the 1st term, d=the common difference
The information provided gives us 2 equations and 2 unknowns:
a(8) = 75 = a + 7d
a(20) = 75 = a + 19d
If both a(8) = 75 and a(20) = 75, then the common difference would be 0, which isn't much of a sequence. You should check the problem again.